• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 08:43:04 on June 9,2007

    Entry number 204874

    keyword=shift summary

    First couple hours spent
    - wait for beam
    - determine settings for 2x2 raster
    - determine rad levels with radiator in
    - wait for beam while Hall B injector work done
    - simultaneously adjust scintillator HV and beta
    - go to holey carbon, play with raster size and beam
    position to find the hole to center the beam -
    cannot get straight through BPMs, settle on slightly
    skewed path
    - get Tiefenback energy set correctly at 3 AM

    4 AM: beam going away for 15 min - 1 hour for work on Hall B

    5 AM: beam back, start looking at gamma's on 3He to see if we understand trigger rate and heating of radiator. Initial slope of ~3C/minute and apparent asymptote of dT ~ 80 C match estimates for
    20 W deposited in target - done at 20 uA current on 6% foils.

    5:30 AM on... taking data gamma 3He -> pp+n to check out system.

    left arm ioc dead.
    we have total of about 30k t5 triggers taken at 20uAmp beam.

    problems for the next shift:

    1. fix focal plan polarimater readout -cable out for 4u
    2. fix problem with lef arm IOC
    3. estabilsh a quality control on line replay